OpenStudy (anonymous):

If n >= 1 is an integer, prove that 2^(2n) - 1 is divisible by 3. Is there a formal way to prove this or can I just demonstrate it by plugging in n = 1,2,3,etc?

OpenStudy (anonymous):

Proving by "plugging in" would involve an infinite number of cases.

OpenStudy (experimentx):

try using induction

OpenStudy (anonymous):

My problem is the only induction I know is plugging examples in!

OpenStudy (anonymous):

OK, how did this happen: 4(2^(2n) - 1 + 1) - 1 = 4 - 1?

OpenStudy (experimentx):

let it put it up again, \[ 2^{2n + 2} - 1 = 4 \cdot2^{2n} - 1 = 4 (2^{2n} - 1 + 1)- 1\] you know 3 divides 2^(2n) - 1 from assumption of induction. \[ 4 (2^{2n} - 1) + 4 - 1 = 4 \cdot (2^{2n }-1) + 3 \] since 3 divides both terms, it divides the whole thing.

OpenStudy (experimentx):

also notice that \[ 2^{2n} = 4^n = (3 + 1)^n = 3 \times (\text{something}) + 1 \] you can use this to show that the reminder is 1

OpenStudy (anonymous):

When I assume what I'm trying to prove to be true, what are the things I must do to make the whole argument valid?

OpenStudy (experimentx):

do you know how to use induction?

OpenStudy (anonymous):

this is my first problem in induction? Semester started this week.

OpenStudy (anonymous):

avoiding induction, I will give another way, \[f(n)= 4^n-1 \] the values of the sum of the digits of f(n) takes a function represented by g(n): \[3n+1 \]thus 3n+1-1 is divisible by 3

OpenStudy (experimentx):

Oh!! first put few values of n to show that it is true like, n=1,2,3 ... up to some number 'k' if it holds for k+1 assuming it holds for k, then it holds for any finite number.

OpenStudy (experimentx):

@Ahmad1 seeing that made me remember \[ 4^n - 1 = (4 - 1)\cdot (1 + 4 + 4^2 + \dots + 4^{n-1})\] it holds trivially

OpenStudy (anonymous):

I don't understand this "the values of the sum of the digits of f(n) takes a function represented by g(n):3n+1"

OpenStudy (anonymous):

@Ahmad1 what did I miss?

OpenStudy (experimentx):

that is just another way of looking at the problem 4^n -1 = 3k 4^n = 3k + 1 for some integer k

OpenStudy (anonymous):

Oh I see. I get it now. Thanks everybody.

OpenStudy (experimentx):

have seen this theorem? http://en.wikipedia.org/wiki/Fermat's_little_theorem

OpenStudy (anonymous):

nope and I don't know modular arithmetic.